Prove that |a+b| - |a-b| ≤ 2|b|.


1
Expert's answer
2022-06-08T17:39:58-0400

Consider paralelogram "ABCD"

"\overrightarrow{AB}=\vec a, \overrightarrow{AD}=\vec b"

Then


"\overrightarrow{AO}=\dfrac{1}{2}(\vec{a}+\vec {b}), \overrightarrow{DO}=\dfrac{1}{2}(\vec{a}-\vec {b})"

The sum of any two sides of a triangle is greater than or equal to the third side.


"\triangle AOD:AD+OD\ge AO"

Then


"|\vec{b}|\ge |\dfrac{1}{2}(\vec{a}+\vec {b})|-|\dfrac{1}{2}(\vec{a}-\vec {b})|"

"|\vec{a}+\vec {b}|-|\vec{a}-\vec {b}|\le 2|\vec{b}|"